# Applying the maximum entropy principle to neural networks enhances multi-species distribution models

### v3

@@ 3.1 Comparative analysis of SDM methods
Table 2 shows the performances of various standard comparative methods, including Maxent, Boosted Regression Tree (BRT) with or without TGB correction, and the neural network model for multi-species proposed by Zbinden et al. (2024), all evaluated with the average AUC per region and the general average  (Phillips et al., 2006; Valavi et al., 2022; Zbinden et al., 2024). It also contains the performance of our main DeepMaxent implementation (best architecture and hyperparameters) and the baseline losses, with or without the TGB correction.
@@
CE 0.701 0.661 0.732 0.724 0.772 0.793 0.731
CE (using TGB) 0.727 0.708 0.739 0.732 0.771 0.792 0.745
BCE 0.656 0.600 0.718 0.736 0.804 0.799 0.719
BCE (using TGB) 0.723 0.726 0.743 0.739 0.803 0.846 0.763
Poisson loss 0.658 0.599 0.714 0.737 0.804 0.799 0.719
Poisson loss (using TGB) 0.712 0.727 0.732 0.731 0.800 0.846 0.758
@@
DeepMaxent 0.654 0.593 0.718 0.744 0.803 0.810 0.720
DeepMaxent (using TGB) 0.714 0.732 0.752 0.754 0.803 0.850 0.767
@@
Without TGB sampling bias correction, performances are close among methods and range from 0.718 to 0.723 in general average AUC (Table 2), except for the CE loss which is much better (0.731). Except for the latter, we observe no general performance gain for the tested deep learning losses (BCE, Poisson, DeepMaxent, ranging from 0.719 to 0.720) compared to the literature methods, e.g. Maxent (0.721) or the best SDM Ensemble of Valavi et al. (2022) (0.723).
@@
The TGB correction brings a consistent general performance improvement for all methods, including the ones of the literature and our implementations. However, not all approaches respond equally strongly to the sampling bias correction. For instance, Maxent gains 0.039 in general averaged AUC by using TGB, and it is the same for BRT. Regarding our implemented baseline losses, TGB induces an AUC gain of 0.011 for the CE loss, 0.044 for the BCE loss and 0.039 for the Poisson loss. Finally, DeepMaxent gains 0.047, making it the best method in terms of general AUC (0.767). These results show that the proposed DeepMaxent is particularly adapted to this bias correction technique while it enables to leverage the predictive potential of neural networks for multi-species spatial intensity estimation. The largest region average AUC gains were mostly seen in regions CAN and AWT, where spatial sampling bias is the strongest according to  (Phillips et al., 2009). Note that the best method of Zbinden et al. (2024), achieving a general averaged AUC of 0.755, incorporated both random and TGB points as absences in their BCE loss, and their results specifically showed the key role of the TGB points in this performance. DeepMaxent-TGB also had the best AUC in four of the six regions (CAN, NSW, NZ, SWI), showing that it is robust across regions and biological groups (NSW includes four biological groups, see Table 1). BCE-TGB is the second-best method in general AUC (0.763). In contrast, CE-TGB and Poisson-TGB yield poorer general AUC (0.745 and 0.758) than Maxent-TGB (0.760) or BRT-TGB (0.759).
@@ 3.2 Sensitivity study
Table 3 shows the average AUCs calculated for all regions, according to six different values for each hyperparameter: batch size, number of hidden layers and weight decay. A detailed analysis of AUC values for each region is provided in the appendix (D). The general performance of DeepMaxent-TGB was quite robust to hyper-parameter choices and neural network depth. In particular, DeepMaxent-TGB kept a general average AUC above 0.764, i.e. above all other methods, for all tested batch sizes, ranging from 10 to 2500. Qualitatively, a smaller batch size induces smoother species intensity maps, while larger batch size tends to concentrate the intensity in higher abundance areas, as illustrated for one species in the region CAN in Figure 2. The L2 regularisation (weight decay) has a larger impact on DeepMaxent-TGB performance. The regularisation has a small but consistently positive impact on the performance up to a value of $3\times 10^{-4}$ (see Table 3), while further increasing the weight decay value results on a performance degradation due to oversmoothing (see Figure 2). DeepMaxent-TGB keeps the highest general average AUC among tested methods for $\tau$ ranging from $3\times 10^{-4}$ to $1\times 10^{-3}$ Varying the number of hidden layers in the neural network architecture of DeepMaxent from one to two had almost no effect, with a same general averaged AUC of 0.767 (Table 3), and the AUC softly and progressively decreased for three (0.766), four (0.764), five (0.762) and six layers (0.759).
Table 3: Average AUC values for DeepMaxent-TGB across all regions, calculated for six different values of each hyperparameter: batch size, number of hidden layers, and weight decay. The default values used are a batch size of 250, two hidden layers, and a weight decay of 3e-4.
Batch size AUC Hidden layers AUC Weight decay AUC
@@
Smaller batch size Lower weight decay
Reference
Larger batch size Higher weight decay
Figure 2: Estimated probabilities for the species can14 (CAN) by varying batch size and weight decay, while keeping other hyperparameters at their default values (batch size = 250, hidden layers = 2, weight decay = $3\times 10^{-4}$) corresponding to an AUC of 0.919. Results are shown (i) for different batch sizes: (i) a lower batch size (10) with AUC of 0.923, (ii) a reduced weight decay ($1\times 10^{-5}$) with an AUC of 0.904, and (iii) a higher batch size (2500) with an AUC of 0.921, while a larger weight decay ($1\times 10^{-1}$) with an AUC of 0.555.
@@ 4 Discussion
The proposed DeepMaxent method, which jointly learns the spatial density for multiple species through a scalable batched algorithm and allows inducing results with maximum entropy, was evaluated on a benchmark consisting of six distinct regional datasets. These include different biological groups and covariates. Compared to state-of-the-art methods, DeepMaxent with the TGB bias correction achieved the highest average AUC across all regions and outperforms all other methods in four of the six regions (see Table 2). It notably surpasses the single-species Maxent and BRT methods with TGB correction, as well as the recent multi-species neural network method of Zbinden et al. (2024). We applied the TGB correction for spatial sampling bias correction with presence-only data Phillips et al. (2009) to DeepMaxent by restricting training sites to the ones including at least one observation of the biological group from which the species are predicted. We showed that DeepMaxent is particularly adapted to the TGB correction as it outperformed the same multi-species model learned with three other commonly used loss functions in this context: The Poisson loss from which it derives (Renner and Warton (2013)), the cross-entropy over species (Deneu et al. (2021); Brun et al. (2024)) and the binary-cross-entropy loss (encoding occurrences as presence and pseudo-absences, Benkendorf and Hawkins (2020); Zbinden et al. (2024)). We believe this is an important result in the sense that it demonstrates that (i) our methodology for extending Maxent to deep learning is functional on biased presence-only data, (ii) it outperforms the original Maxent method on the same input tabular data and (iii) it outperforms alternative loss functions used for multi-species SDM based on neural networks. Further efforts could adapt different background sampling strategies while respecting dataset specificity for biological groups and species (Schartel and Cao, 2024).
@@
For feature extraction, a fully-connected neural network with just one or two hidden layers was found to be optimal for these datasets which comprise only tabular data. This simple MLP architecture performs well on tasks involving low-dimensional and unstructured data. While deep learning is often associated with large models, simpler architectures can be more efficient and beneficial for such tasks. For multi-species settings, DeepMaxent reduces the overall computational cost by learning a single feature extractor shared across species (Ba and Caruana, 2014; Raghu et al., 2017), notably compared to methods relying on more resource-intensive, per-species, calculations (Merow et al., 2014).
@@
More broadly, the proposed approach is flexible regarding the type of input and species observation data and should facilitate data integration approaches in the future by using neural networks. In the case of more structured input data, the neural network architecture could be adapted to ingest different types of structured inputs (e.g., sequential or spatial), which might lead to capturing complementary spatio-temporal environmental patterns (Deneu et al. (2021); Estopinan et al. (2022)). Importantly, DeepMaxent is the first method to bridge the gap between deep learning and point process-based SDM (Renner et al. (2015)). Using the formulation of the Poisson process within the loss function, our method could seamlessly incorporate additional loss terms to integrate diverse and more standardised types of species observations, such as presence-absence surveys (Fithian et al. (2015)), detection/non-detection histories (Koshkina et al. (2017)), and abundance or imperfect count data (Dorazio (2014)). This approach, called integrated SDMs, have recently been highlighted as a promising avenue to enhance the reliability of SDMs (Miller et al. (2019); Isaac et al. (2020); Mostert and O’Hara (2023)). Extending DeepMaxent with this approach could use standardised datasets to disentangle the real abundance of each species from detection biases, while harnessing the extensive geographic coverage of opportunistic presence-only data.
@@
The DeepMaxent loss was notably more adapted to the TGB bias correction technique compared to the BCE, CE and Poisson losses. Regarding BCE with TGB, which was the second-best method overall, it suggests that interpreting TG background sites as absences is reasonable as it filters many false absences, but not optimal as it gives too much importance to the remaining false absences. An apparent paradox is that, given the equivalence between the Maxent and Poisson process losses (Renner and Warton (2013)), which applies to any model for the spatial intensity (e.g. neural networks) and induces the equivalence between the DeepMaxent-TGB and Poisson-TGB losses, we would expect similar results for the two latter losses, but they were different. We see two hypothetical algorithmic explanations that could jointly play a role in this discrepancy: (i) the intensity scale in the Poisson loss, which cancels in the DeepMaxent loss, affects the gradient on the other parameters determining the density, the learning trajectory and final estimates, and (ii) the stochastic batched algorithm interacts with (i) in deviating the two estimates. Even though the CE loss had the best results without TGB correction, and thus appears natively less sensitive to spatial sampling bias, its performances with TGB remained below the best methods which relied on this correction (DeepMaxent, BCE, Poisson, Maxent). This limited performance may be due to the loss of information on the spatial variations of the intensity for each species when normalising the intensity across species for each site. More broadly, regarding the estimation of multiple species spatial intensities from biased presence-only data, it suggests that learning to classify the most likely observed species per site is a limited approach.
Additionally to the smoothing of species spatial densities brought by the loss function of DeepMaxent, which induces entropy maximisation, this study also highlighted that the L2 regularisation induced more spatial smoothing which enhances model performance in a deep learning setting. This spatial smoothing can be related to the effect of the L2 regularisation of Maxent Phillips et al. (2006) and, more broadly, to regularised loss functions improving the robustness of point process based SDMs Komori et al. (2024). However, excessively high weight decay values ($\tau$) can be detrimental, resulting penalising model weights too harshly. This can lead to a over-smoothed species spatial densities which don’t capture the actual spatial variations of a species abundance.
@@
We provided a mathematical guarantee to justify the use of a stochastic batched gradient descent algorithm to learn DeepMaxent on batches of sites, and we further showed that varying the batch size (from 10 to 2500) had a relatively weak impact on general performances. From a computational perspective, such algorithm drastically reduces the need for computer memory when training DeepMaxent compared to optimising the loss function on all sites at each iteration, as only the data for one batch needs to be loaded at a time in memory. Yet, the batch size may affect the learning trajectory due to the approximation of the partition function, and we noticed its influence on the final model performance. Similarly to the increasing the weight decay, we observed that decreasing the batch size may smooth species spatial densities. Although batch size is one of the less sensitive hyperparameters, identifying an optimal batch size remains important and it may interact with e.g. the learning rate and number of epochs. The use of the algorithm Adam Kingma (2014), known for its robustness to hyperparameter adjustments, contributed to the stability of DeepMaxent performances.
One classical limitation of PO-based SDMs is the fact that, in many cases, no PA data is available to validate the model hyperparameters. Although we have seen that DeepMaxent is rather robust to hyperparameter choices, including mini-batch size, number of hidden layers, and weight decay value, finding the right hyperparameters is required to obtain optimal results. In order to mitigate this limitation, we verified that hyper-parameter tuning by validating the results on PO data results on good values being chosen, thus removing the need for PA data during model training. These results can be found in the Appendix C.
Acknowledgements
Funded by the European Union. B3 (Biodiversity Building Blocks for policy) receives funding from the European Union’s Horizon Europe Research and Innovation Programme (ID No 101059592).

### v4

@@ 3.1 Comparative analysis of SDM methods
Table 2 shows the performances of various traditional SDM methods, including Maxent, Boosted Regression Tree (BRT) with or without TGB correction, and the recent neural network model for multi-species proposed by Zbinden et al. (2024), all evaluated with the average AUC per region, along with overall average (Phillips et al., 2006; Valavi et al., 2022; Zbinden et al., 2024). It also contains the performance of our main DeepMaxent implementation and the baseline deep learning losses, with and without the TGB correction.
@@
CE 0.701 0.661 0.732 0.724 0.772 0.793 0.731 $\pm$ 0.001
CE (using TGB) 0.727 0.708 0.739 0.732 0.771 0.792 0.745 $\pm$ 0.001
BCE 0.656 0.600 0.718 0.736 0.804 0.799 0.719 $\pm$ 0.002
BCE (using TGB) 0.722 0.730 0.743 0.738 0.804 0.849 0.764 $\pm$ 0.002
Poisson loss 0.658 0.599 0.714 0.737 0.804 0.799 0.719 $\pm$ 0.002
Poisson loss (using TGB) 0.712 0.730 0.732 0.729 0.801 0.849 0.759 $\pm$ 0.002
@@
DeepMaxent 0.654 0.593 0.718 0.744 0.803 0.810 0.720 $\pm$ 0.001
DeepMaxent (using TGB) 0.712 0.732 0.752 0.753 0.806 0.850 0.768 $\pm$ 0.001
@@
Without TGB sampling bias correction, performances are overall lower and close among methods, ranging from 0.716 to 0.723 in overall average AUC (Table 2), except for the CE loss which achieves 0.731. Except for the latter, we observe no general performance gain for the tested deep learning losses (BCE, Poisson, DeepMaxent, ranging from 0.719 to 0.720) compared to the literature methods, e.g. Maxent (0.721) or the best SDM Ensemble of Valavi et al. (2022) (0.723).
@@
The TGB bias correction brings a consistent performance improvement for all methods, including the ones in the literature and our implementations. However, not all approaches respond equally strongly to TGB. For instance, Maxent gains 0.039 in overall averaged AUC by using TGB, and it is the same for BRT. Regarding our implemented baseline losses, TGB induces an AUC gain of 0.011 for the CE loss, 0.045 for the BCE loss and 0.040 for the Poisson loss. Finally, DeepMaxent achieves an improvement of 0.048 with TGB, resulting in the highest overall AUC (0.768). These results show that the proposed DeepMaxent is well adapted to this bias correction technique while it enables to leverage the predictive potential of multi-species neural networks for spatial density estimation. The largest region average AUC gains were mostly seen in regions CAN and AWT, where spatial sampling bias is the strongest according to (Phillips et al., 2009). Note that the best method of Zbinden et al. (2024), achieving an overall averaged AUC of 0.755, incorporated both random and TGB points as absences in their BCE loss, and their results specifically showed the key role of the TGB points in this performance. DeepMaxent with TGB also had the best AUC in four of the six regions (CAN, NSW, NZ, SWI), showing that it is robust across regions and biological groups (NSW includes four biological groups, see Table 1). BCE with TGB is the second-best method in overall AUC (0.764). In contrast, using TGB, CE and Poisson yield poorer overall AUC (0.745 and 0.759) than Maxent (0.760) or BRT (0.759).
Indeed, although we would expect similar results for DeepMaxent and Poisson when using the same TGB strategy, given the equivalence of their global minimizer (see Appendix A.1, DeepMaxent resulted in a 0.023 gain in AUC, closing the gap towards a perfect score by close to 10% (i.e. 10% less mis-ordered pairs of presence and absence sites) compared to Poisson, a significant gain according to statistical tests (Appendix E).
Figure 2 shows AUC scores by observation abundance class (rare, common and abundant) for each loss function, averaged over all regions. Averaging takes into account region-specific abundance class distributions (see appendix 21).
Figure 2: Comparison of average AUC values across all regions by loss and abundance classes on NCEAS dataset
DeepMaxent consistently outperforms the other losses across all abundance classes, while BCE ranks second in every class. Poisson loss performs moderately well for abundant and common species. For rare species, the Poisson loss is dominated by the many zero-observation terms, which directly penalize the predicted intensities via the $\lambda_{ij}$ component. This tends to enforce uniformly low intensity values, potentially leading to underfitting beyond what data scarcity alone would justify. In contrast, CE shows the opposite trend: it achieves its best results for rare species, but lower AUC scores for abundant and common species, suggesting that its sensitivity to class imbalance limits its overall effectiveness in these abundance regimes. These results suggest that DeepMaxent is particularly well-suited for modelling species with heterogeneous distributions and low occurrence counts, where classical Poisson-based methods may underperform due to their tendency to over-penalize predictions in data-sparse regions.
(a) PO data for can01
(b) PA data for can01
(c) DeepMaxent - AUC of 0.942
(d) BCE - AUC of 0.926
(e) CE - AUC of 0.932
(f) Poisson - AUC of 0.923
Figure 3: Estimated relative probabilities for the species can01 (a rare species with 16 PO points): (a) Presence-Only (PO) data, yellow points; (b) Presence-Absence (PA) data where green corresponds to presences, and magenta to absences. (c–f) Estimated from different loss functions: (c) DeepMaxent; (d) Binary Cross-Entropy (BCE); (e) Cross-Entropy (CE); (f) Poisson loss.
Estimated probabilities for can01, a rare species, differ widely depending on the loss function applied (see Fig. 3). For this species, presence-only (PO) data are rare, while presence-absence (PA) data are more abundant. In contrast, the maps generated for can02, a common species, give much more consistent estimates between the different methods (see appendix 12).
Table 3 reports the general averaged AUC values for different loss functions evaluated on two datasets, Bioclim-GeoPlant and Landsat-GeoPlant, using TGB sampling bias correction.
Table 3: Comparison of method performance by general averaged AUC. The best average AUC for each column is highlighted in bold, while the second-best averaged AUC is underlined. All methods use TGB for the background samples.
Loss AUC
Results for Bioclim-GeoPlant
Maxent 0.823
CE 0.830 $\pm$ 0.001
BCE 0.839 $\pm$ 0.001
Poisson loss 0.837 $\pm$ 0.002
DeepMaxent 0.860 $\pm$ 0.001
Results for Landsat-GeoPlant
CE 0.829 $\pm$ 0.001
BCE 0.885 $\pm$ 0.002
Poisson loss 0.862 $\pm$ 0.002
DeepMaxent 0.887 $\pm$ 0.001
On the Bioclim-GeoPlant dataset, the DeepMaxent model achieves the highest average AUC (0.860), outperforming all losses, including BCE (0.839), Poisson (0.837), and CE (0.830), as well as the Maxent model (0.823). These results highlight the improved predictive capacity of DeepMaxent when using environmental variables from Bioclim. A similar result is observed on the Landsat-GeoPlant dataset, where DeepMaxent also performs best with a general averaged AUC of 0.887, closely followed by BCE (0.885), while CE and Poisson yield lower performances (0.829 and 0.862, respectively). DeepMaxent performs well on satellite time series data. BCE Loss also performs well in this context, especially for a large dataset. Notably, the standard deviations across random initializations are very small for all models ($\leq$0.002), indicating high consistency across runs.
@@ 3.2 Sensitivity study
Table 4 shows the average AUCs calculated for all regions, according to six different values for each hyper-parameter: mini-batch size, number of hidden layers and weight decay. A detailed analysis of AUC values for each region is provided in the appendix (D). The general performance of DeepMaxent-TGB was quite robust to hyper-parameter choices, with the largest difference in average AUC across all tested values being only 0.010, illustrating the model’s stability with respect to mini-batch size, number of hidden layers, and weight decay. In particular, DeepMaxent with TGB kept a general average AUC above 0.764, i.e. above the best results using all other methods, for all tested mini-batch sizes, ranging from 10 to 2500. Qualitatively, a smaller mini-batch size induces smoother species intensity maps, while larger mini-batch size tends to concentrate the intensity in higher abundance areas, as illustrated for one species in the region CAN in Figure 4. The L2 regularisation (weight decay) has an important impact on DeepMaxent performance. It has a small but consistently positive impact on the performance up to a value of $3\times 10^{-4}$ (see Table 4), while further increasing the weight decay value results on a performance degradation due to oversmoothing (see Figure 4). Varying the number of hidden layers in the neural network architecture of DeepMaxent from one to two had almost no effect, with a same general averaged AUC of 0.767 (Table 4), and the AUC softly and progressively decreased for three (0.766), four (0.764), five (0.762) and six layers (0.759).
Table 4: Average AUC values for DeepMaxent-TGB across all regions, calculated for six different values of each hyper-parameter: mini-batch size, number of hidden layers, and weight decay. The default values used are a mini-batch size of 250, two hidden layers, and a weight decay of 3e-4.
Mini-batch size AUC Hidden layers AUC Weight decay AUC
@@
Smaller value Higher value
can01
Batch size
Weight decay
can02
Batch size
Weight decay
Figure 4: Estimated relative probabilities for the species can01 and can02 (CAN) by varying mini-batch size and weight decay, while keeping other hyper-parameters at their default values (mini-batch size = 250, hidden layers = 2, weight decay = $3\times 10^{-4}$). Smaller mini-batch sizes correspond to 10, and larger sizes to 3000. For weight decay, smaller values correspond to $3\times 10^{-6}$, and larger values to 0.1.
@@ 4 Discussion
In this study, we propose DeepMaxent, a new SDM method for PO data that generalizes Maxent to multi-species deep neural networks with a direct relationship to a Poisson count loss. DeepMaxent can be trained with a scalable batched algorithm, and implicitly implements the Target-Group Background sampling bias correction (TGB) initially proposed for Maxent (Phillips et al., 2009). We conduct an extensive evaluation of the method on two SDM benchmark datasets: NCEAS (Elith et al., 2020) and GeoPlant (Picek et al., 2025), which together span a wide range of species, biological groups, regions, and data regimes. The NCEAS dataset enables comparison with a broad set of state-of-the-art SDM methods previously evaluated on this dataset (Phillips et al., 2009; Valavi et al., 2022; Zbinden et al., 2024), such as Maxent and BRT. We also use NCEAS to compare the DeepMaxent loss against batched implementations of alternative losses (Poisson, BCE, CE), and to assess the efficiency of the implicit TGB bias correction compared to uniform random background for each loss. Consistent with previous findings on traditional SDM methods (Phillips et al., 2009; Ranc et al., 2017; Barber et al., 2022), deep learning-based SDMs benefit strongly and systematically from TGB regardless of the loss function. DeepMaxent is well adapted to the TGB correction as it outperforms the same model architecture trained using the three alternative losses: the Poisson loss from which it derives, the cross-entropy over species (Deneu et al., 2021; Brun et al., 2024)) and the binary-cross-entropy loss (encoding occurrences as presence and pseudo-absences(Benkendorf and Hawkins, 2020; Zbinden et al., 2024)). DeepMaxent with TGB achieves the highest average AUC across all NCEAS regions and outperforms all other methods in four of the six regions (see Table 2). It notably surpasses the single-species Maxent and BRT methods with TGB correction.
More broadly, our results illustrate that, like other SDM methods, deep learning-based SDMs must properly account for spatial sampling biases to reveal their performance potential. This can be achieved through the compatibility of DeepMaxent and TGB, but many other proposed bias correction methods could bring further improvements to DeepMaxent (e.g. Boria et al. (2014); Botella et al. (2021)). Future efforts could consider more sophisticated bias correction strategies, such as adapting background sampling strategies to biases specific to each biological group and species, as these biases are often a major source of variation in the performance of species distribution models (Schartel and Cao, 2024), or explicitly model spatial sampling effort (Warton et al., 2013; Botella et al., 2021).
@@
For feature extraction on the NCEAS dataset, DeepMaxent with a fully-connected neural network with just two hidden layers is found to be optimal across most regions. This simple MLP architecture performs well on tasks involving low-dimensional and unstructured tabular data. While deep learning is often associated with large models, simpler architectures can be more efficient and beneficial for such tasks. For more complex tasks such as the GeoPlant dataset, DeepMaxent is compatible with any type of neural network architecture, allowing the model to obtain the best overall results on Landsat-GeoPlant using a convolutional architecture for time series.
In multi-species settings, DeepMaxent, like other deep learning–based SDMs (and some specific approaches such as concurrent ordination (van der Veen et al., 2023)), reduces overall computational cost by learning a single feature extractor shared across species (Ba and Caruana, 2014; Raghu et al., 2017), in contrast to most methods that rely on more resource-intensive, per-species computations (Merow et al., 2014). Similar ideas of reducing multi-species complexity into shared latent dimensions exist in statistical ecology (e.g. concurrent ordination (van der Veen et al., 2023) and community-level driver models (Ovaskainen et al., 2017a)), but DeepMaxent implements it through deep neural feature learning rather than explicit latent-variable modelling. It is important to note that each layer added increases the number of model parameters. Although these parameters are shared between the different species, this increase in complexity can lead to overfitting, particularly when data are limited or noisy. A prudent approach, therefore, is to start with simple architectures and then gradually make the model more complex, by adding hidden layers, as long as validation performance continues to improve. A deterioration in validation performance then serves as a warning signal that the model is becoming too complex in relation to the amount of information available in the training data.
@@
More broadly, the proposed approach is flexible regarding the type of input and species observation data and should facilitate data integration approaches in the future by using neural networks. In the case of more structured input data, the neural network architecture could also be adapted to ingest other types of inputs (e.g. spatial remote sensing imagery), which might lead to capturing complementary spatio-temporal environmental patterns (Deneu et al., 2021; Estopinan et al., 2022). Importantly, DeepMaxent is the first method to bridge the gap between deep learning and point process-based SDM (Renner et al., 2015). Using a point-process formulation of species distributions within our loss framework, the present work focuses on presence-only data through a Poisson likelihood. However, this formulation is general and could, in principle, be extended to incorporate other types of ecological observations, such as presence–absence surveys (Fithian et al., 2015), detection/non-detection histories (Koshkina et al., 2017), or abundance and imperfect count data (Dorazio, 2014). In such extensions, all observation types would remain linked to the same underlying predicted ecological intensity, while each type would be associated with its own appropriate observation likelihood. When these likelihoods involve additional parameters (e.g., detection probabilities), they could be jointly estimated within the same optimisation framework or through hierarchical extensions under suitable independence assumptions (see, e.g., Isaac et al. (2020)).
These approaches for combining various observation types, called integrated SDMs, have recently been highlighted as a promising avenue to enhance the reliability of SDMs (Miller et al., 2019; Isaac et al., 2020; Mostert and O’Hara, 2023). Extending DeepMaxent with this approach could use standardised datasets to disentangle the real relative abundance of each species from detection biases, while harnessing the extensive geographic coverage of opportunistic presence-only data. To further increase ecological realism, the DeepMaxent framework could be extended to incorporate species co-occurrence patterns, following the principles of Joint Species Distribution Models (Pollock et al., 2014). This integration, which involves modelling the joint probability of species occurrences given the environment, has already proven feasible within deep learning-based SDM architectures (Chen et al., 2017).
Maxent and log-linear Poisson regression have been shown to be equivalent in terms of the estimated probability across sites (Renner and Warton, 2013), and we found a similar equivalence between the more general DeepMaxent and Poisson losses, in that they share the same global minimizer when applied to a single species (Appendix A.1). This would lead us to expect similar results for DeepMaxent and Poisson, but DeepMaxent consistently outperformed Poisson in our experiments. One key difference that could explain the discrepancy lies in the treatment of the absolute intensities in the multi-species setting: in the un-normalised Poisson loss, the model will try to capture differences in observation counts between species, leading to more importance in the loss for frequently observed species. In contrast, the DeepMaxent loss normalises intensities across sites for each species, which removes the effect of the number of observations in the per-species loss magnitude and decouples parameter updates from absolute intensity. Additionally, the stochastic mini-batch optimisation interacts with this normalisation, further differentiating the parameter estimates between the two losses.
@@
Even though the CE loss has the best results without TGB correction, and thus appears natively less sensitive to spatial sampling bias, its performances with TGB remain below the rest of methods, which benefit more from this correction (DeepMaxent, BCE, Poisson, Maxent). This limited performance may be due to the loss of information on the spatial variations of the intensity for each species when normalising the intensity across species for each site. More broadly, regarding the estimation of multiple species spatial intensities from biased presence-only data, these results suggest that learning to classify the most likely observed species (Deneu et al., 2021; Estopinan et al., 2022; Brun et al., 2024) per site leads to sub-optimal results.
In addition to the smoothing of species spatial densities achieved by the DeepMaxent loss function, which induces entropy maximization, this study highlights that L2 regularisation can be used to further encourage smoothing of predicted probabilities (see Figure 4). By penalizing large model parameters, L2 reduces overfitting and produces smoother predictions, resulting in more gradual changes in species probability across space and improved performance in presence-only settings, potentially alleviating overfitting in the low-data regime.
@@
Concerning the optimisation process of DeepMaxent, we provided a mathematical guarantee to justify the use of a stochastic batched gradient descent algorithm and we further showed that varying the batch size (from 10 to 2500) had little impact on general performances (see Table 4). From a computational perspective, this algorithm is highly scalable for large datasets, as it drastically reduces memory and computational requirements by processing only one mini-batch at a time. Moreover, it can take full advantage of GPU parallelization, further accelerating training and enabling efficient handling of high-dimensional or large-scale data (see H). Yet, the mini-batch size may affect the learning trajectory due to the approximation of the partition function, and we noticed its influence on the final model behaviour. Similarly to increasing the weight decay, we observed that decreasing the mini-batch size may smooth species spatial densities. Although mini-batch size is one of the least sensitive hyper-parameters, identifying an optimal mini-batch size remains important, and it may interact with other hyper-parameters such as the learning rate and number of epochs.
One classical limitation of PO-based SDMs is that presence–absence (PA) data are often unavailable to validate model hyper-parameters. To address this issue, we evaluated hyper-parameter tuning directly on PO data and found that it reliably identifies suitable values, thereby removing the need for PA data during model training. The corresponding analyses are provided in the Appendix C.
Overall, DeepMaxent provides improvements over both traditional presence-only SDMs and previous deep learning-based approaches on a variety of case studies. It leverages neural networks to learn complex, high-order non-linear relationships directly from data, enabling flexible and expressive modelling, while benefiting from the principles behind the success of the original Maxent. However, several methodological directions remain open. Future extensions could integrate additional ecological information, such as species traits or phylogenetic relationships, to model inter-species dependencies more explicitly, an approach explored in models like HMSC (Ovaskainen et al., 2017a). Additionally, DeepMaxent could offer a framework for developing and testing bias-reduction strategies in deep learning-based SDMs, such as improved background sampling or observation models, to enhance robustness in heterogeneous and biased datasets.